# Quantum-vacuum-induced topological edge polaritons

### v1

@@ Abstract
This paper uncovers the formation of topological edge polaritons that are induced by the presence of quantum vacuum. Such quantum-vacuum-induced edge polaritons could be achieved in a system of spinful fermionic lattice under appropriate interaction with a single photonic mode. In the absence of the light-matter coupling, the system is shown to be topologically trivial, which consequently does not support edge modes. By employing Floquet theory, the system is also found to be topologically trivial in the classical light limit, i.e., at very small light-matter coupling but very large number of photons. On the other hand, by treating both the fermionic and photonic degrees of freedom quantum mechanically, the system becomes topologically nontrivial in the full (fermionic+photonic) Hilbert space, which manifests itself as a pair of topological (almost) zero energy eigenstates localized near each lattice’s edge and with very small mean photon number. Finally, the robustness of such quantum-vacuum-induced edge polaritons against spatial disorder and counterrotating coupling effect is explicitly demonstrated.
@@ I Introduction
Quantum vacuum has remained a central subject in the area of quantum electrodynamics that underlies various physical phenomena such as the Casimir effect Casimir48 , spontaneous emission Frisch33 , the Lamb shift Lamb47 , and (very recently) quantum-vacuum-induced superconductivity Schlawin19 . In particular, these phenomena arise due to the interaction between particles and virtual photons, which are induced by quantum vacuum fluctuations that result from the presence of such particles. Observing the quantum effects of light, including quantum vacuum, involves access to the regime of strong light-matter coupling in systems of particles that interact with a sufficiently low number of photons. This is typically achieved by placing such particles inside an optical cavity/quantum resonators, which provides a controllable platform for trapping particles and photons, as well as for engineering appropriate interaction among them. Such a physical setup has been extensively studied under the timely area of cavity quantum electrodynamics Vidal21 , and its implementation in the lab has been successfully demonstrated Landig16 ; Chikkaraddy16 ; Santhosh16 .
@@
This work uncovers another mechanism under which quantum vacuum yields a physical observable with no classical counterpart. Specifically, by noticing that quantum vacuum mimics a physical edge of a one-dimensional (1D) system, it is capable of supporting the analogues of zero energy edge modes (as typically found in 1D topological insulators Asboth16 and topological superconductors Kit ), under an appropriately constructed Hamiltonian. It is worth noting that zero energy edge modes are particularly attractive for quantum technological applications. Indeed, a pair of the so-called Majorana zero modes that emerge at the ends of 1D topological superconductors collectively form a nontrivial qubit, which consequently exhibits natural protection against physical errors Nayak08 . For this reason, efforts towards realizing Majorana-based qubits in experiments Lutchyn10 ; Oreg10 ; Mourik12 ; Nichele17 ; Das12 ; Rokhinson12 ; Lee14 and harnessing their quantum computational power Karzig17 ; Litinski17 ; Landau16 ; Plugge16 ; Litinski18 ; Viyuela19 have been extensively made since over the last two decades and are still ongoing Paetznick23 . Non-Majorana zero modes, such as those that emerge at the ends of 1D topological insulators or their bosonic variations, also find applications in other aspects of quantum technologies despite their lack in nonlocal properties. For example, such modes were recently utilized as additional degrees of freedom to encode quantum information locally at one end of a topological quantum circuit Mei2018 . By adiabatically tuning appropriate system parameters, the encoded quantum information could then be transmitted to the other end of the circuit with a high fidelity Mei2018 , thus achieving the process of quantum state transfer.
@@
The ability of light to induce topological features in an otherwise trivial system may be reminiscent of Floquet topological phases Lindner11 ; Jiang11 . However, it is worth noting that the mechanism underlying our proposed system in this paper is fundamentally different from that underlying Floquet topological phases. In particular, Floquet topological phases arise when light interacts classically with the system at hand Lindner11 ; Jiang11 , thus rendering its Hamiltonian time-periodic. As a result of this time-periodicity, energy is no longer conserved and is replaced by the so-called quasienergy, which is only defined modulo the driving frequency Shirley65 ; Sambe73 . Due to the periodicity of the quasienergy space, additional band touching events may arise and consequently induce a topological phase transition, thus resulting in a variety of novel topological phases, often with no static counterparts Bomantara16 ; Rudner13 ; Nathan15 ; Asboth14 ; Ho14 ; Bomantara20 ; Zhou18 ; Zhu22 ; Bomantara22 ; Koor22 . Apart from the emergence of band touching events at the edges of the quasienergy Brillouin Zone above, Floquet topological phases may also arise when the (classical) light-matter interaction renormalizes some system parameters in the effective static Hamiltonian description, thus rendering the system topologically nontrivial even when its undriven counterpart is topologically trivial Ino10 ; Kitagawa10 ; Oka09 ; Calvo11 ; Kitagawa11 ; Morell12 . By contrast, the topological effect arising in the present work is not caused by any of the aforementioned mechanisms, but instead results from the fact that the photonic Fock space is bounded from below by the quantum vacuum. Indeed, without taking into account the quantum effect of light, such a topological feature will not be present.
@@
Finally, it is worth noting that studies of Floquet topological phases under full quantum light treatment have developed into an active research area since the last decade Trif12 ; Gulacsi15 ; Bomantara16b ; Ciuti21 ; Chio21 ; Dmytruk22 ; Mendez20 ; Winter23 ; Bacciconi24 ; Karzig15 . However, to our knowledge, the role of quantum vacuum in manifesting topological effects has remained elusive. In particular, while the topology of quantum vacuum itself has been studied in the context of high energy physics Volovik13 ; Alexander20 , its implication on a physical observable has yet to be identified. This work aims to close this knowledge gap by explicitly demonstrating the presence of topological edge modes in a system with strong light-matter coupling at very low photon mean numbers. This is achieved by devising an appropriate 1D topologically trivial fermionic lattice that interacts with a single photonic mode via a Tavis-Cummings-like coupling. Such an interaction renders the composite light-matter system (polariton) topologically nontrivial in the enlarged Hilbert space containing the fermionic and the photonic degrees of freedom. Remarkably, as the quantum vacuum state behaves as an “edge” in the photonic Fock space, the system topology manifests itself as in-gap states that are only present near the lattice’s (physical) edges and at very low photon mean numbers. That is, such edge states are not observable in the classical light limit and their existence is linked to the presence of quantum vacuum, as intended.
@@
This manuscript is organized as follows. In Sec. II.1, we mathematically present the Hamiltonian describing the proposed system. In Sec. II.2, we elucidate the intuition behind the system’s construction by explicitly demonstrating the emergence of quantum-vacuum-induced topological edge polaritons at special parameter values. In Sec. II.3, we numerically calculate the system’s energy spectrum and identify the quantum-vacuum-induced topological edge polaritons at more general parameter values. There, we further verify that the quantum-vacuum-induced topological edge polaritons have an almost zero mean photon number. In Sec. III, we verify the absence of any topological edge mode in the system when the light-matter coupling is treated classically. In Sec. IV, we demonstrate the robustness of our quantum-vacuum-induced topological edge polaritons against spatial disorder and counterrotating coupling term in the system’s Hamiltonian. Finally, we summarize our results and discuss prospects for future studies in Sec. V.
II Topological edge modes induced by quantum vacuum
@@ II.1 Model description
We consider a 1D chain of spinful fermions interacting with a photon mode under the Hamiltonian ($\hbar=1$ units are used throughout this paper)
$H=H_{\rm fer}+H_{\rm ph}+H_{\rm int}\;,$ (1)
@@
$H_{\rm fer}$ $=$ $\sum_{s=\pm 1}\left\{\sum_{j=1}^{N}J_{1}c_{s,A,j}^{\dagger}c_{s,B,j}+\sum_{j=1}^{N-1}J_{2}c_{s,B,j}^{\dagger}c_{s,A,j+1}\right.$
@@
$c_{s,X,j}$ ($c_{s,X,j}^{\dagger}$) is the fermionic annihilation (creation) operator with spin $s=\pm\frac{1}{2}$, sublattice $X\in\left\{A,B\right\}$, and at the $j$th site of $N$ lattice sites, $a$ ($a^{\dagger}$) is the bosonic annihilation (creation) operator, $\omega$ is the corresponding photon energy, $J_{1}$ and $J_{2}$ are respectively the intra- and inter-site hopping amplitudes, $E_{Z}$ is the magnetic field strength, and $\Delta$ is the photon-fermion interaction strength.
$H_{\rm fer}$ describes a pair of the Su-Schrieffer-Heeger (SSH) models Su1980 , each of which is known to support a topological phase. Indeed, in the special case of $J_{1}=E_{Z}=0$, the edge fermionic operators $c_{\pm,A,1}$ and $c_{\pm,B,N}$ commute with $H_{\rm fer}$, consequently representing its zero modes. At $J_{1}\neq 0$ but with $E_{Z}=0$, zero modes remain present as long as $J_{1}<J_{2}$, though they no longer correspond to $c_{\pm,A,1}$ and $c_{\pm,B,N}$ anymore. It should be noted, however, that such zero modes cease to exist as $E_{Z}$ becomes nonzero. In particular, the magnetic field term of $H_{\rm fer}$ hybridizes the zero modes localized on the same edge but associated with different spin sectors, thereby lifting their degeneracy. Therefore, the Hamiltonian $H_{\rm fer}$ as a whole is not topologically protected, despite it being made up of two topological models.
$H_{\rm ph}$ describes the photonic part of the Hamiltonian, whilst $H_{\rm int}$ describes the interaction part that couples the fermionic and photonic degrees of freedom. The interaction term could be interpreted as a modified Tavis-Cummings (TC) term Tavis68 which involves a fermion hopping from one sublattice to the other in addition to flipping its spin state. Note that in potential experimental realizations of our model, the spin degree of freedom could be equivalently replaced by a pseudospin one, e.g., the ground and excited states of a two-level superconducting qubit Fink09 . In this work, we assume the framework of a single fermionic particle occupying our 1D lattice, which in turn significantly reduces the complexity of our model. While this eliminates the observation of interesting phenomena typically associated with the TC model such as superradiance Tavis68 , it suffices to demonstrate the topological effects induced by quantum vacuum.
@@ II.2 Intuitive insight at special parameter values
$\mathcal{S}=\sum_{j=1}^{N}\sum_{s=\pm 1}\sum_{X=A,B}sc_{s,X,j}^{\dagger}c_{s,X,j}+a^{\dagger}a$ (3)
then serves as a conserved quantity with respect to the full Hamiltonian $H$. Consequently, the set of states $\left\{|+,n,X,j\rangle,|-,n+1,X,j\rangle:X\in\left\{A,B\right\},j=1,\cdots,N\right\}$, where $c_{\pm,X,j}^{\dagger}c_{\pm,X,j}|\pm,n,X,j\rangle=|\pm,n,X,j\rangle$ and $a^{\dagger}a|\pm,n,X,j\rangle=n\omega|\pm,n,X,j\rangle$, is closed under the application of $H$.
@@
Figure 1(a) shows the schematics of our Hamiltonian under the special case $J_{1}=E_{Z}=0$ presented above, from which two important observations are made. First, the $(-,n+1)$ and $(+,n)$ spin-photon sectors coupled by the Hamiltonian can be deformed into two independent SSH models with an effective intra-(inter-)site coupling strength of $\Delta\sqrt{n+1}$ ($J_{2}$). Consequently, a pair of zero modes exist at each lattice edge for spin-photon sectors that satisfy $\Delta\sqrt{n+1}<J_{2}$. In particular, if $\Delta>J_{2}$, zero modes do not exist for any pair of $(-,n+1)$ and $(+,n)$ spin-photon sectors. Second, note that the $(-,0)$ spin-photon sector is detached from the other sectors and behaves as an SSH model in the topologically nontrivial regime regardless of the photon-fermion interaction strength. As a result, at the special parameter values under consideration, the full Hamiltonian $H$ always supports a pair of zero modes in the $(-,0)$ spin-photon sector, which correspond to the states $|-,0,A,1\rangle$ and $|-,0,B,N\rangle$. Since such edge modes live in the Hilbert space spanned by the combined fermionic and photonic degrees of freedom, we refer to them as quantum-vacuum-induced edge polaritons. In the following, we will discuss the fate of these quantum-vacuum-induced edge polaritons in cases for which either $E_{Z}$ or $J_{1}$ is non-zero.
Case i: $E_{Z}\neq 0$ and $J_{1}=0$
@@
In this case, the $(-,0)$ spin-photon sector is no longer decoupled. However, by additionally setting $\omega=0$ for simplicity, we find that a pair of zero modes still exist, but now correspond to the superposition states which can be found up to first order in $J_{2}$ as (see Appendix A for details)

### v2

# Quantum-vacuum-protected topological edge polaritons

@@ Abstract
This paper uncovers the formation of topological edge polaritons that are protected by the presence of quantum vacuum. Such quantum-vacuum-protected edge polaritons could be achieved in a system of neutral atomic lattice under appropriate interaction with a single photonic mode. In the absence of the light-matter coupling, the system is shown to be topologically trivial, which consequently does not support edge modes. By employing Floquet theory, the system is also found to be topologically trivial in the classical light limit, i.e., at very small light-matter coupling but very large number of photons. On the other hand, by treating both the atomic and photonic degrees of freedom quantum mechanically, the system becomes topologically nontrivial in the full (atomic+photonic) Hilbert space, which manifests itself as a pair of topological (almost) zero energy eigenstates localized near each lattice’s edge and with very small mean photon number. Finally, the robustness of such quantum-vacuum-protected edge polaritons against spatial disorder and counterrotating coupling effect is explicitly demonstrated.
@@ I Introduction
Quantum vacuum has remained a central subject in the area of quantum electrodynamics that underlies various physical phenomena such as the Casimir effect Casimir48 , spontaneous emission Frisch33 , the Lamb shift Lamb47 , and (very recently) quantum-vacuum-induced superconductivity Schlawin19 . In particular, these phenomena arise due to the interaction between particles and virtual photons, which are induced by quantum vacuum fluctuations that result from the presence of such particles. Observing the quantum effects of light, including quantum vacuum, involves access to the regime of strong light-matter coupling in systems of particles that interact with a sufficiently low number of photons. This is typically achieved by placing such particles inside an optical cavity/quantum resonators, which provides a controllable platform for trapping particles and photons, as well as for engineering appropriate interaction among them. Such a physical setup has been extensively studied under the timely area of cavity quantum electrodynamics Sentef18 ; Vidal21 ; Schlawin22 ; Torre21 , and its implementation in the lab has been successfully demonstrated Landig16 ; Chikkaraddy16 ; Santhosh16 .
@@
This work uncovers another mechanism under which quantum vacuum yields a physical observable with no classical counterpart. Specifically, by noticing that quantum vacuum mimics a physical edge of a one-dimensional (1D) system, it is capable of supporting the analogues of zero energy edge modes (as typically found in 1D topological insulators Asboth16 and topological superconductors Kit ), under an appropriately constructed Hamiltonian. It is worth noting that zero energy edge modes are particularly attractive for quantum technological applications. Indeed, a pair of the so-called Majorana zero modes that emerge at the ends of 1D topological superconductors collectively form a nontrivial qubit, which consequently exhibits natural protection against physical errors Nayak08 . For this reason, efforts towards realizing Majorana-based qubits in experiments Lutchyn10 ; Oreg10 ; Mourik12 ; Nichele17 ; Das12 ; Rokhinson12 ; Lee14 and harnessing their quantum computational power Karzig17 ; Litinski17 ; Landau16 ; Plugge16 ; Litinski18 ; Viyuela19 have been extensively made since over the last two decades and are still ongoing Paetznick23 . Non-Majorana zero modes, such as those that emerge at the ends of 1D topological insulators or their bosonic variations, also find applications in other aspects of quantum technologies despite their lack in nonlocal properties. For example, such modes were recently utilized as additional degrees of freedom to encode quantum information locally at one end of a topological quantum circuit Mei2018 . By adiabatically tuning appropriate system parameters, the encoded quantum information could then be transmitted to the other end of the circuit with high fidelity Mei2018 , thus achieving the process of quantum state transfer.
@@
The ability of light to induce topological features in an otherwise trivial system may be reminiscent of Floquet topological phases Lindner11 ; Jiang11 . However, it is worth noting that the mechanism underlying our proposed system in this paper is fundamentally different from that underlying Floquet topological phases. In particular, Floquet topological phases arise when light interacts classically with the system at hand Lindner11 ; Jiang11 , thus rendering its Hamiltonian time-periodic. As a result of this time-periodicity, energy is no longer conserved and is replaced by the so-called quasienergy, which is only defined modulo the driving frequency Shirley65 ; Sambe73 . Due to the periodicity of the quasienergy space, additional band touching events may arise and consequently induce a topological phase transition, thus resulting in a variety of novel topological phases, often with no static counterparts Bomantara16 ; Rudner13 ; Nathan15 ; Asboth14 ; Ho14 ; Bomantara20 ; Zhou18 ; Zhu22 ; Bomantara22 ; Koor22 . Apart from the emergence of band touching events at the edges of the quasienergy Brillouin Zone, Floquet topological phases may also arise when the (classical) light-matter interaction renormalizes some system parameters in the effective static Hamiltonian description, thus rendering the system topologically nontrivial even when its undriven counterpart is topologically trivial Ino10 ; Kitagawa10 ; Oka09 ; Calvo11 ; Kitagawa11 ; Morell12 . By contrast, the topological effect arising in the present work is not caused by any of the aforementioned mechanisms, but instead results from the fact that the photonic Fock space is bounded from below by the quantum vacuum. Indeed, without taking into account the quantum effect of light, such a topological feature will not be present.
@@
Finally, it is worth noting that studies of Floquet topological phases under full quantum light treatment have developed into an active research area since the last decade Trif12 ; Gulacsi15 ; Bomantara16b ; Ciuti21 ; Chio21 ; Dmytruk22 ; Mendez20 ; Winter23 ; Bacciconi24 ; Karzig15 ; Wang19 . However, to our knowledge, the role of quantum vacuum in manifesting topological effects has remained elusive. In particular, while the topology of quantum vacuum itself has been studied in the context of high energy physics Volovik13 ; Alexander20 , its implication on a physical observable has yet to be identified. This work aims to close this knowledge gap by explicitly demonstrating the presence of topological edge modes in a system with strong light-matter coupling at very low photon mean numbers. This is achieved by devising an appropriate 1D topologically trivial lattice of neutral atoms or artificial atoms that interacts with a single photonic mode via a Tavis-Cummings-like coupling, which could arise in cold atom or superconducting circuit systems. Such an interaction renders the composite light-matter system (polariton) topologically nontrivial in the enlarged Hilbert space containing the atomic and the photonic degrees of freedom. Remarkably, as the quantum vacuum state behaves as an “edge” in the photonic Fock space, the system topology manifests itself as in-gap states that are only present near the lattice’s (physical) edges and at very low photon mean numbers. That is, such edge states are not observable in the classical light limit and their existence is linked to the presence of quantum vacuum, as intended. It should however be emphasized that we take a broad definition of quantum vacuum in this paper, which does not necessarily correspond to a true electromagnetic vacuum. The notion of quantum vacuum used in this paper may generally refer to the zero number state that serves as the boundary of any effective Fock space.
@@
This manuscript is organized as follows. In Sec. II.1, we mathematically present the Hamiltonian describing the proposed system. In Sec. II.2, we elucidate the intuition behind the system’s construction by explicitly demonstrating the emergence of quantum-vacuum-protected topological edge polaritons at special parameter values. In Sec. II.3, we numerically calculate the system’s energy spectrum and identify the quantum-vacuum-protected topological edge polaritons at more general parameter values. There, we further verify that the quantum-vacuum-protected topological edge polaritons have an almost zero mean photon number. In Sec. III, we verify the absence of any topological edge mode in the system when the light-matter coupling is treated classically. In Sec. IV, we demonstrate the robustness of our quantum-vacuum-protected topological edge polaritons against spatial disorder and counterrotating coupling term in the system’s Hamiltonian. Finally, we summarize our results and discuss prospects for future studies in Sec. V.
II Topological edge modes protected by quantum vacuum
@@ II.1 Model description
We consider a 1D chain of two-level neutral atoms or artifical atoms interacting with a photon mode under the Hamiltonian ($\hbar=1$ units are used throughout this paper)
$H=H_{\rm atom}+H_{\rm ph}+H_{\rm int}\;,$ (1)
@@
$H_{\rm atom}$ $=$ $\sum_{s=\pm 1}\left\{\sum_{j=1}^{N}J_{1}c_{s,A,j}^{\dagger}c_{s,B,j}+\sum_{j=1}^{N-1}J_{2}c_{s,B,j}^{\dagger}c_{s,A,j+1}\right.$
@@
$c_{s,X,j}$ ($c_{s,X,j}^{\dagger}$) is the particle annihilation (creation) operator with its two internal levels labeled as $s=\pm 1$, sublattice $X\in\left\{A,B\right\}$, and at the $j$th site of $N$ lattice sites, $a$ ($a^{\dagger}$) is the bosonic annihilation (creation) operator, $\omega$ is the corresponding photon energy, $J_{1}$ and $J_{2}$ are respectively the intra- and inter-site hopping amplitudes, $E_{Z}$ is the (artificial) magnetic field strength, and $\Delta$ is the light-matter interaction strength.
$H_{\rm atom}$ describes a pair of the Su-Schrieffer-Heeger (SSH) models Su1980 , each of which is known to support a topological phase. Indeed, in the special case of $J_{1}=E_{Z}=0$, the edge operators $c_{\pm,A,1}$ and $c_{\pm,B,N}$ commute with $H_{\rm atom}$, consequently representing its zero modes. At $J_{1}\neq 0$ but with $E_{Z}=0$, zero modes remain present as long as $J_{1}<J_{2}$, though they no longer correspond to $c_{\pm,A,1}$ and $c_{\pm,B,N}$ anymore. It should be noted, however, that such zero modes cease to exist as $E_{Z}$ becomes nonzero. In particular, the magnetic field term of $H_{\rm atom}$ hybridizes the zero modes localized on the same edge but associated with different internal levels, thereby lifting their degeneracy. Therefore, the Hamiltonian $H_{\rm atom}$ as a whole is not topologically protected, despite it being made up of two topological models.
$H_{\rm ph}$ describes the photonic part of the Hamiltonian, whilst $H_{\rm int}$ describes the interaction part that couples the atomic and photonic degrees of freedom. The interaction term could be interpreted as a modified Tavis-Cummings (TC) term Tavis68 which involves an atom hopping from one sublattice to the other in addition to flipping its internal level state. Note that in potential experimental realizations of our model, the internal levels could be represented by the ground and excited states of a two-level superconducting qubit Fink09 . In this work, we assume the framework of a single neutral particle occupying our 1D lattice, which in turn significantly reduces the complexity of our model. While this eliminates the observation of interesting phenomena typically associated with the TC model such as superradiance Tavis68 , it suffices to demonstrate the topological effects protected by quantum vacuum.
@@ II.2 Intuitive insight at special parameter values
$\mathcal{S}=\sum_{j=1}^{N}\sum_{s=\pm 1}\sum_{X=A,B}\frac{s}{2}c_{s,X,j}^{\dagger}c_{s,X,j}+a^{\dagger}a$ (3)
then serves as a conserved quantity with respect to the full Hamiltonian $H$ (the eigenvalue $\ell=n+s/2$ of $\mathcal{S}$ is a good quantum number). Consequently, the set of states $\left\{|+,n,X,j\rangle,|-,n+1,X,j\rangle:X\in\left\{A,B\right\},j=1,\cdots,N\right\}$, where $c_{\pm,X,j}^{\dagger}c_{\pm,X,j}|\pm,n,X,j\rangle=|\pm,n,X,j\rangle$ and $a^{\dagger}a|\pm,n,X,j\rangle=n|\pm,n,X,j\rangle$, is closed under the application of $H$.
@@
Figure 1(a) shows the schematics of our Hamiltonian under the special case $J_{1}=E_{Z}=0$ presented above, from which two important observations are made. First, the $(-,n+1)$ and $(+,n)$ atom-photon sectors coupled by the Hamiltonian can be deformed into two independent SSH models with an effective intra-(inter-)site coupling strength of $\Delta\sqrt{n+1}$ ($J_{2}$) and with an effective sublattice dependent onsite potential that is $\propto\omega$. Consequently, a pair of edge modes might exist at each lattice edge for atom-photon sectors that satisfy $\Delta\sqrt{n+1}<J_{2}$. Nevertheless, such edge modes are not topological even at these special parameter values due to the presence of the onsite potential. Moreover, for $\Delta\gg J_{2}$, such edge modes are absent for any pair of $(-,n+1)$ and $(+,n)$ atom-photon sectors. Second, note that the $(-,0)$ atom-photon sector is detached from the other sectors and behaves as an SSH model in the topologically nontrivial regime regardless of the light-matter interaction strength. As a result, at the special parameter values under consideration, the full Hamiltonian $H$ always supports a pair of zero modes in the $(-,0)$ atom-photon sector, which correspond to the states $|-,0,A,1\rangle$ and $|-,0,B,N\rangle$.
Figure 1(b) shows the schematics of our Hamiltonian when all system parameters are nonzero. In this case, the system is no longer separated into decoupled sectors and instead forms an effective 2D lattice spanned by the combined atomic and photonic degrees of freedom, in which all sites are nontrivially connected. This effective 2D lattice has a qualitatively similar structure to a typical second-order topological insulator STI1 ; STI2 ; Bomantara19 ; STI3 ; STI4 ; STI5 ; STI6 ; STI7 , which is known to support topologically protected corner modes. However, due to the the dependence of the coupling in the “photon number direction” on $\sqrt{n+1}$, the effective 2D Hamiltonian of our system is not exactly the same as that of any existing second-order topological insulator, and its full topological characterization is much more difficult. Nevertheless, upon identifying the presence of corner modes under the special parameter values above, the system’s second-order topology could still be established by verifying that such corner modes are indeed robust at more general parameter values as detailed in the following. To gain further intuition along the way, we will start by analyzing the simplest generalization by setting $E_{Z}=0$ and $J_{1}\neq 0$, under which the system is still separated into decoupled sectors. We will then consider the slightly more complicated case of $E_{Z}\neq 0$ and $J_{1}=0$, under which all sites are coupled and the full effective 2D lattice must be accounted for. Finally, the most general case of all system parameters being nonzero will be discussed in Sec. II.3.
Case i: $E_{Z}=0$ and $J_{1}\neq 0$
In this case, $\ell=n+s/2$ remains a good quantum number with which the Hamiltonian can be divided into decoupled sectors via
where $\langle\ell|=\left(\langle-,0,X,j|\right)$ for $\ell=-1/2$ and $\langle\ell|=\left(\langle+,n,X,j|,\langle-,n+1,X,j|\right)$ for $\ell=n+1/2$, $X=A,B$, and $j=1,2,\cdots,N$. The effective $2\times 2$ Hamiltonian describing each decoupled sector $\ell$ can be written as
$\mathcal{H}_{\rm eff,\ell=-1/2}$ $=$ $\left[J_{1}+J_{2}\cos(\hat{k})\right]\sigma_{x}+J_{2}\sin(\hat{k})\sigma_{y},$
$\mathcal{H}_{\rm eff,\ell\neq-1/2}$ $=$ $\left[J_{1}+J_{2}\cos(\hat{k})\right]\sigma_{x}+J_{2}\sin(\hat{k})\sigma_{y}$ (5)
$+\sqrt{\ell+1/2}\Delta\tau_{x}\sigma_{x}+\frac{\omega}{2}\tau_{z},$
where we have suppressed unimportant constant terms, $\hat{k}$ is the quasimomentum operator such that $e^{\pm\mathrm{i}\hat{k}}$ shifts the lattice site ($j$) by one unit, $\sigma$’s and $\tau$’s are respectively the Pauli matrices acting on the sublattice ($A,B$) and $-\ell$ ($(+,n),(-,n+1)$) degrees of freedom.
$\mathcal{H}_{\rm eff,\ell=-1/2}$, which corresponds to the $(-,0)$ atom-photon sector, remains mathematically identical to the paradigmatic SSH model and thus continues to support a pair of zero modes as long as $J_{1}<J_{2}$. The remaining pairs of $(-,n+1)$ and $(+,n)$ atom-photon sectors are effectively described by $\mathcal{H}_{\rm eff,\ell\neq-1/2}$, which also satisfies the chiral symmetry $\mathcal{C}^{\dagger}\mathcal{H}_{\rm eff,\ell}(\hat{k})\mathcal{C}=-\mathcal{H}_{\rm eff,\ell}(\hat{k})$, particle-hole symmetry $\mathcal{P}^{\dagger}\mathcal{H}_{\rm eff,\ell}(\hat{k})\mathcal{P}=-\mathcal{H}_{\rm eff,\ell}(-\hat{k})$, and time-reversal symmetry $\mathcal{T}^{\dagger}\mathcal{H}_{\rm eff,\ell}(\hat{k})\mathcal{T}=\mathcal{H}_{\rm eff,\ell}(-\hat{k})$, but under different operators $\mathcal{C}=\tau_{x}\sigma_{z}$, $\mathcal{P}=\mathcal{K}\tau_{x}\sigma_{z}$, and $\mathcal{T}=\mathcal{K}$ respectively, where $\mathcal{K}$ is the complex conjugation operator. Therefore, all $\mathcal{H}_{\rm eff,\ell}$ belongs to the BDI class according to the Altland-Zirnbauer classification AZ97 , the topology of which is characterized by a $\mathbb{Z}$ invariant.
where
which is defined under the assumption of periodic boundary conditions (PBC) such that $\hat{k}\rightarrow k\in[-\pi,\pi)$ is a good quantum number. We numerically compute the winding number $w_{+}$ at four different $\ell$ sectors in Fig. 2. In particular, by varying $J_{2}$ while fixing $J_{1}=\Delta=1$, Fig. 2(a) reveals that $w_{+}$ is trivial everywhere, indicating the absence of a topological phase transition at $J_{1}=J_{2}$ in all these nonzero photon sectors. Figure 2(b) further demonstrates that $w_{+}$ is also 00 for all $\Delta$ values under consideration, even in the regime $J_{1}<J_{2}$ where each SSH model is topologically nontrivial. While not shown in the figure, the same results are also observed for $w_{-}$. Our winding number calculations thus confirm the absence of topological edge modes at $\ell\neq-1/2$. Therefore, under $E_{Z}=0$, the system only supports a pair of topological edge modes in the quantum vacuum sector of the Hamiltonian, i.e., as described by $\mathcal{H}_{\rm eff,\ell=-1/2}$.
Figure 2: The winding number $w_{+}$ (a) as a function of $J_{2}$ at a fixed $\Delta=J_{1}=1$ and $\omega=2\pi$, and (b) as a function of $\Delta$ at a fixed $J_{2}=0.5$, $J_{1}=0.1$, and $\omega=2\pi$.
Case ii: $E_{Z}\neq 0$ and $J_{1}=0$
@@
In this case, the $(-,0)$ atom-photon sector is no longer decoupled. However, by additionally setting $\omega=0$ for simplicity, we find that a pair of zero modes still exist, but now correspond to the superposition states which can be found up to first order in $J_{2}$ as (see Appendix A for details)